OpenStudy (anonymous):

how can ^100 be simplified

OpenStudy (anonymous):

what does ^100 mean?

OpenStudy (anonymous):

2^100=(2^2)^50=4^50=(4^2)^25=16^25=(16^5)^5=1048576^5 which is some big number

jhonyy9 (jhonyy9):

like (*10)*10

jhonyy9 (jhonyy9):

or exponent 2 on exponent 50

jhonyy9 (jhonyy9):

orexp 25 on exp 4

jhonyy9 (jhonyy9):

or invers exp 4 on exp 25

jhonyy9 (jhonyy9):

in every ways how you can getting the result 100 from multiplication